OpenStudy (tw101):

Point X is located at (2, -6) and point Z is located at (0, 5). Find the y value for the point Y that is located the distance from point X to point Z. (4 points) -3.6 -3.8 -4 -4.2

OpenStudy (anonymous):

Wait, hm, is that the correct graph for this problem? The question mentions X Y and Z and I can't seem to find any of those on the shape.

OpenStudy (tw101):

oh whoops! sorry wrong equation haha ignore that

OpenStudy (anonymous):

It's ok! :P

OpenStudy (tw101):

okay so you were saying

OpenStudy (tw101):

okay so you were saying

OpenStudy (tw101):

The bottom question is what i was asking

OpenStudy (anonymous):

Is the original question the one you are confused on? the point X is located... etc?

OpenStudy (karatechopper):

Is Y somewhere between X and Z? Almost forming a triangle shaped...?

OpenStudy (tw101):

sorry about that one sec

OpenStudy (tw101):

I figured out the weird triangle one, now I just have a question about the point x one

OpenStudy (karatechopper):

Please rewrite the question below my comment, I am a bit confused about which exact question. Is the one that is your current question?

OpenStudy (tw101):

Point X is located at (2, -6) and point Z is located at (0, 5). Find the y value for the point Y that is located the distance from point X to point Z. (4 points) -3.6 -3.8 -4 -4.2

OpenStudy (anonymous):

@tw101 , are you familiar with the 'Distance Formula' at all?

OpenStudy (abdullahm):

I believe it wants the midpoint...

OpenStudy (anonymous):

It says that point Y is located at the distance of point X to point Z, not between them I believe. But I could be mistaken.

OpenStudy (tw101):

isnt it c^2=a^2+b^2?

OpenStudy (abdullahm):

"Find the y value for the point Y that is located the distance from point X to point Z." It could be asking for what y-value is the same distance from x as from z.

OpenStudy (abdullahm):

Also, trying to find the distance of it would not get us anywhere and midpoint makes more sense.

OpenStudy (anonymous):

Ah, okay that makes sense now

OpenStudy (abdullahm):

\(\sf\Large Midpoint=(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\) and that is for two points: \(\sf (x_1,y_1)\) and \(\sf (x_2,y_2)\)

OpenStudy (tw101):

okay give me a second to work it out:)

OpenStudy (abdullahm):

mhmm, the answer isn't there. So it isn't the midpoint :/

OpenStudy (tw101):

wait im confused what do you mean

OpenStudy (tw101):

i didnt get the right answer either...

OpenStudy (tw101):

so what method can i use then..?

OpenStudy (abdullahm):

I'm not sure :/

OpenStudy (tw101):

dang.. okay well thank you for your help anyways.